OpenStudy (awy):

Find the vaule of: lim x approaches 0: (x^2+x)/(2*sqrt(x^3+x^2))

OpenStudy (raden):

factor out the numerator and then denominator. x^2 + x = x(x+1) and 2*sqrt(x^3+x^2) = 2 sqrt(x^2(x+1)) = 2x sqrt(x+1) then you can cancel the x's

OpenStudy (tkhunny):

Do we know if we are approaching from the right or from the left? If we don't then \(\sqrt{x^{2}} = |x|\). It makes quite a difference when playing around zero (0).

OpenStudy (awy):

This helped, thank you.

OpenStudy (awy):

the question didnt state left or right.

OpenStudy (awy):

just lim as x approaches 0.

OpenStudy (tkhunny):

With no clarification from the problem statement, you must consider each case separately. |x| = x for \(x \ge 0\) |x| = -x for \(x < 0\)